# Haoran Guan

### v1

@@ 5 Numerical experiments
5.1 The settings and background of our experiments
We show the specification of our computer below in Table 1.
Table 1: The specification of our computer
@@ 5.1 The settings and background of our experiments
In this section, we conduct numerical experiments using a matrix derived from lower-triangular matrices. Such matrices frequently arise in the fields of numerical PDEs, financial mathematics, and the solution of linear systems (see Practical , Theory , Geometric , Elements , and their additional references). We construct the input matrix $X$ by stacking several small square matrices from top to bottom. Specifically, we utilize a small lower-triangular matrix $X_{1}\in\mathbb{R}^{n\times n}$, defined as follows:
@@
$x_{1ij}$ $=$ $0,i=1,2,\cdots,n,j>i,.$
@@
Here, $a$ is a negative constant. $X_{1}$ is a matrix which occurs in some real problems in the industry. We form $X\in\mathbb{R}^{m\times n}$ based on $b$ small $X_{1}$, that is, $X=[X_{1}^{\top},X_{1}^{\top},\cdots,X_{1}^{\top}]^{\top}$. $b$ is a positive integer and $b=\frac{m}{n}$. In the following group of experiments, we very $a$, $b$ and $n$ to do numerical experiments to show the properties of our algorithms. If $X_{ss}=\Omega_{2}\Omega_{1}X$, $X\in\mathbb{R}^{m\times n}$, we let $\epsilon_{1}=0.5$, $p_{1}=0.6$, $\epsilon_{2}=0.5$ and $p_{2}=0.4$ in Definition 2. We take different $m$, $n$, $s_{1}$ and $s_{2}$ in the following experiments. Therefore, $\epsilon_{b}=1.25$, $\epsilon_{s}=0.75$. We can easily find that Assumption 1 is satisfied if $mn$ do not exceed the level of $10^{12}$. This is enough for the problems in the engineering. Regarding the single-sketching, we take the Gaussian sketch. Here, we have $\epsilon=0.5$ and $p=0.6$. When $mn$ do not exceed the level of $10^{12}$, our settings can also meet the requirement in Assumption 2. Readers are encouraged to conduct experiments with other types of sketches as well.
5.2 Tests of the applicability and the numerical stability of the algorithms
@@
In this group of experiments, we focus on the applicability and numerical stability of our algorithms, comparing them with the original LU-CholeskyQR2. First, we examine the potential influence of $\kappa_{2}(X)$. We conduct two groups of numerical experiments with different values of $m$: $m=20000$ and $m=30000$. We fix $n=50$ and vary $a$ from $-70$, $-80$, $-90$ to $-100$ to alter $\kappa_{2}(X)$. For the sizes of $\Omega_{1}$ and $\Omega_{2}$, we set $s_{1}=17000$ and $s_{2}=50$ for SLHC2, and $s=50$ for SSLHC3, based on (9) and (10). Our numerical experiments are summarized in Tables 24, and Table 6. Next, we shift our focus to $n$. We fix $a=70$ and vary $n$ from $5$, $10$, $20$ to $50$ while conducting numerical experiments. Here, we set $s_{1}=\frac{n^{2}+n}{0.5\cdot 0.6^{2}}=\frac{n^{2}+n}{0.15}$ and $s_{2}=50$ for SSLHC3, while $s=50$ for SLHC2. We perform two groups of numerical experiments with $m=20000$ and $m=30000$. The results of these experiments are listed in Table 6 and Table 7.
@@
According to Tables 27, we find that our LHC2, SLHC2, and SSLHC3 demonstrate significant advantages over the original LU-CholeskyQR2 in terms of applicability when $X$ is ill-conditioned. This is a key innovative point of our work. Among the three new algorithms, our deterministic LHC2 exhibits good numerical stability in both orthogonality and residuals. However, the orthogonality of SLHC2 is less stable and relatively worse than that of LHC2 and SSLHC3. This instability arises because SLHC2 lacks the CholeskyQR2 structure that underpins SLHC. This can also be observed in the comparisons between (25), (99), and (106). In the randomized algorithms, the orthogonality and residuals of SSLHC3 are nearly on par with those of the deterministic LU-CholeskyQR2 and LHC2. This demonstrates the effectiveness of multi-sketching in ensuring numerical stability. Our SSLHC3 effectively balances numerical stability and applicability.
Table 2: Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
@@
Table 3: Comparison of residual with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
@@
Table 4: Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
@@
Table 5: Comparison of residual with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
@@
Table 6: Comparison of residual with different $n$ varying when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 7: Comparison of residual with different $n$ varying when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
5.3 Tests of CPU times of the algorithms
@@
Table 8: Comparison of CPU time(s) between different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 9: Comparison of CPU time(s) between different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 10: Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=20000$ and $n=20$
@@
Table 11: Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=30000$ and $n=20$
@@
Table 12: Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 13: Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
Table 14: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 15: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 16: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 17: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
Table 18: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 19: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 20: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 21: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
5.4 Tests of the robustness of the randomized algorithms
In this section, we present a series of distinct numerical experiments that, to our knowledge, have not been conducted in other similar works. We will assess the robustness of our two randomized algorithms after performing a certain number of experiments. Specifically, we will examine the influence of $\kappa_{2}(X)$, the size of the sketching matrix, and $n$ on the robustness of the algorithms. We conduct two groups of experiments with $m=20000$ and $m=30000$ in each case. To test the effect of different $\kappa_{2}(X)$, we fix $n=50$. For SLHC2, we set $s=50$. For SSLHC3, we set $s_{1}=17000$ and $s_{2}=50$. We vary $a$ from $-70$, $-80$, $-90$ to $-100$ to explore the impact on $\kappa_{2}(X)$. We record the number of successful outcomes every $30$ trials after conducting several groups and calculate the average. The numerical results are presented in Table 22 and Table 23. Regarding the size of the sketching matrix, we focus on the Gaussian sketch, which corresponds to $s$ in SLHC2 and $s_{2}$ in SSLHC3. We fix $n=50$ and $a=-100$, varying the size of the Gaussian sketch from $50$, $100$, $200$ to $400$. The average numbers of successful results in $30$ trials after several groups of experiments are listed in Table 24 and Table 25. When considering different values of $n$, we fix $a=-100$. For SLHC2, we maintain $s=50$. For SSLHC3, we fix $s_{2}=50$ and set $s_{1}=\frac{n^{2}+n}{0.15}$. We vary $n$ from $5$, $10$, $20$ to $50$. The numerical results are detailed in Table 26 and Table 27.
@@ 5.4 Tests of the robustness of the randomized algorithms
Table 22: Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=20000$ and $n=50$
@@
Table 23: Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=30000$ and $n=50$
@@
Table 24: Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=20000$ and $n=50$
@@
Table 25: Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=30000$ and $n=50$
@@
Table 26: Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 27: Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$

### v2

@@ 5. Numerical experiments
5.1. The settings and background of our experiments
We show the specifications of our computer below in Table 1.
Table 1. The specifications of our computer
@@ 5.1. The settings and background of our experiments
In this section, we conduct numerical experiments using a matrix derived from lower-triangular matrices. Such matrices frequently arise in the fields of numerical PDEs, financial mathematics, and the solution of linear systems, see [3, 28, 30, 35], and their additional references. We construct the input matrix $X$ by stacking several small square matrices from top to bottom. Specifically, we utilize a small lower-triangular matrix $X_{1}\in\mathbb{R}^{n\times n}$, defined as follows:
@@
$x_{1ij}$ $=$ $0,i=1,2,\cdots,n,j>i.$
Here, $a$ is a negative constant. We have
$X_{1}=\begin{pmatrix}100&\\ a&100&\\ a&a&100\\ \vdots&\vdots&\ddots&\ddots\\ a&a&\cdots&a&100&\\ a&a&\cdots&a&a&100\end{pmatrix}.$
$X_{1}$ is a matrix which occurs in some real problems in the industry. We form $X\in\mathbb{R}^{m\times n}$ based on several small $X_{1}$ and
$X=\begin{pmatrix}X_{s}\\ X_{s}\\ \vdots\\ X_{s}\end{pmatrix}.$
@@
In the following group of experiments, we very $a$ and $n$ to do numerical experiments to show the properties of our algorithms. If $X_{ss}=\Omega_{2}\Omega_{1}X$, $X\in\mathbb{R}^{m\times n}$, we let $\epsilon_{1}=0.5$, $p_{1}=0.6$, $\epsilon_{2}=0.5$ and $p_{2}=0.4$ in Definition 2.8. We take different $m$, $n$, $s_{1}$ and $s_{2}$ in the following experiments. Therefore, $\epsilon_{b}=1.25$, $\epsilon_{s}=0.75$. We can easily find that Assumption 1 is satisfied if $mn$ do not exceed the level of $10^{12}$. This is enough for the problems in the engineering. Regarding the single-sketching, we take the Gaussian sketch. Here, we have $\epsilon=0.5$ and $p=0.6$. When $mn$ do not exceed the level of $10^{12}$, our settings can also meet the requirement in Assumption 2. Readers are encouraged to conduct experiments with other types of sketches as well.
5.2. Tests of the applicability and the numerical stability of the algorithms
@@
In this group of experiments, we focus on the applicability and numerical stability of our algorithms, comparing them with the original LU-CholeskyQR2. First, we examine the potential influence of $\kappa_{2}(X)$. We conduct two groups of numerical experiments with different values of $m$: $m=20000$ and $m=30000$. We fix $n=50$ and vary $a$ from $-70$, $-80$, $-90$ to $-100$ to alter $\kappa_{2}(X)$. For the sizes of $\Omega_{1}$ and $\Omega_{2}$, we set $s_{1}=17000$ and $s_{2}=50$ for SLHC2, and $s=50$ for SSLHC3, based on (2.9) and (2.10). Our numerical experiments are summarized in Tables 24, and Table 6. Next, we shift our focus to $n$. We fix $a=70$ and vary $n$ from $5$, $10$, $20$ to $50$ while conducting numerical experiments. Here, we set $s_{1}=\frac{n^{2}+n}{0.5\cdot 0.6^{2}}=\frac{n^{2}+n}{0.15}$ and $s_{2}=50$ for SSLHC3, while $s=50$ for SLHC2. We perform two groups of numerical experiments with $m=20000$ and $m=30000$. The results of these experiments are listed in Table 6 and Table 7.
@@
According to Tables 27, we find that our LHC2, SLHC2, and SSLHC3 demonstrate significant advantages over the original LU-CholeskyQR2 in terms of applicability when $X$ is ill-conditioned. This is a key innovative point of our work. Among the three new algorithms, our deterministic LHC2 exhibits good numerical stability in both orthogonality and residuals. However, the orthogonality of SLHC2 is less stable and relatively worse than that of LHC2 and SSLHC3. This instability arises because SLHC2 lacks the CholeskyQR2 structure that underpins SLHC. This can also be observed in the comparisons between (3.9), (4.15), and (4.22). In the randomized algorithms, the orthogonality and residuals of SSLHC3 are nearly on par with those of the deterministic LU-CholeskyQR2 and LHC2. This demonstrates the effectiveness of multi-sketching in ensuring numerical stability. Our SSLHC3 effectively balances numerical stability and applicability.
Table 2. Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
@@
Table 3. Comparison of residual with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
@@
Table 4. Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
@@
Table 5. Comparison of residual with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
@@
Table 6. Comparison of residual with different $n$ varying when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 7. Comparison of residual with different $n$ varying when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
5.3. Tests of CPU times of the algorithms
@@
Table 8. Comparison of CPU time(s) between different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 9. Comparison of CPU time(s) between different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 10. Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=20000$ and $n=20$
@@
Table 11. Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=30000$ and $n=20$
@@
Table 12. Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 13. Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
Table 14. Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 15. Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 16. Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 17. Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
Table 18. Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 19. Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 20. Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
@@
Table 21. Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
@@
5.4. Tests of the robustness of the randomized algorithms
In this section, we assess the robustness of our two randomized algorithms after performing a certain number of experiments. Specifically, we will examine the influence of $\kappa_{2}(X)$, the size of the sketching matrix, and $n$ on the robustness of the algorithms. We conduct two groups of experiments with $m=20000$ and $m=30000$ in each case. To test the effect of different $\kappa_{2}(X)$, we fix $n=50$. For SLHC2, we set $s=50$. For SSLHC3, we set $s_{1}=17000$ and $s_{2}=50$. We vary $a$ from $-70$, $-80$, $-90$ to $-100$ to explore the impact on $\kappa_{2}(X)$. We record the number of successful outcomes every $30$ trials after conducting several groups and calculate the average. The numerical results are presented in Table 22 and Table 23. Regarding the size of the sketching matrix, we focus on the Gaussian sketch, which corresponds to $s$ in SLHC2 and $s_{2}$ in SSLHC3. We fix $n=50$ and $a=-100$, varying the size of the Gaussian sketch from $50$, $100$, $200$ to $400$. The average numbers of successful results in $30$ trials after several groups of experiments are listed in Table 24 and Table 25. When considering different values of $n$, we fix $a=-100$. For SLHC2, we maintain $s=50$. For SSLHC3, we fix $s_{2}=50$ and set $s_{1}=\frac{n^{2}+n}{0.15}$. We vary $n$ from $5$, $10$, $20$ to $50$. The numerical results are detailed in Table 26 and Table 27.
@@ 5.4. Tests of the robustness of the randomized algorithms
Table 22. Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=20000$ and $n=50$
@@
Table 23. Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=30000$ and $n=50$
@@
Table 24. Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=20000$ and $n=50$
@@
Table 25. Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=30000$ and $n=50$
@@
Table 26. Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
@@
Table 27. Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$